Hi, I'm trying to understand the topic of Degree of polynomials but I don't
quite get it, and I'm stuck in this problem, how can I determine the degree of
this polynomial?
8x3 + 6x2y + 4xy3 - 2y4

Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

For polynomials in two or more variables, the degree of a term is the sum of the exponents of the variables in the term; the degree of the polynomial is again the maximum of the degrees of all terms in the polynomial. For example, the polynomial x^2y^2 + 3x^3 + 4y has degree 4, the same degree as the term x^2y^2.
However, a polynomial in variables x and y, is a polynomial in x with coefficients which are polynomials in y, and also a polynomial in y with coefficients which are polynomials in x.
x^2y^2 + 3x^3 + 4y = (3)x^3 + (y^2)x^2 + (4y) = (x^2)y^2 + (4)y + (3x^3)
This polynomial has degree 3 in x and degree 2 in y.
-----------------
Your Problem:
8x^3 + 6x^2y + 4xy^3 - 2y^4 has degree 4
As a polynomial in y it has degree 4
As a polynomial in x it has degree 3
==============
Cheers,
Stan H.


Answer by Edwin McCravy(20054)

Hi, I'm trying to understand the topic of Degree of polynomials but Idon't
quite get it, and I'm stuck in this problem, how can I determine the degree of
this polynomial?

8x3 + 6x2y + 4xy3 - 2y4
The degree of a polynomial is the most number of variables
that are multiplied together in any term of the polynomial
So let's write the eponentials as multiplications of variables:
8x3 + 6x2y + 4xy3 - 2y4 becomes
8xxx + 6xxy + 4xyyy - 2yyyy
The longest strings of multiplied letters are xyyy and yyyy.
Both are strings of 4 multiplied letters, so the degree of the
polynomial is 4.
Edwin